August 23, 2007 - Cosmic Telescope
Today was a catch up day. A few points were retaken from the experiments on the 17th. These new points were averaged into the graphs on that page.
Today it was decided that 1250 Volts was a good operating range based on the previous data. The Scintillators have now both been set at a High Voltage of 1250. The best operating range of the thresholds of the scintillators has been determined to be between 225 mV and 275 mV. Today, a zeroing of these best threshold begins. For this series of experiments, the threshold of the bottom will be set at 225, 250, and 275 mV. At each of these three points, the top will be set at 225, 250, and 275 mV. This data on this experiment is on the page of August 24 because it is then that it was finished.